OpenStudy (anonymous):

The length of a rectangle is 3m less than twice its width, and the area of the rectangle is 27m^2 . Find the dimensions of the rectangle.

OpenStudy (anonymous):

area of rectangle = w*l in your question: l = 2w-3 -> length equals twice width less 3 substitute in area = w*l area = w*(2w-3) = 27 w^2 = 27 + 3 w^2 = 30 w = sqrt(30) if w = sqrt(30) from area = w*l = 27 l = 27 / sqrt(30)